OpenStudy (anonymous):

State how many imaginary and real zeros the function has. f(x) = x3 + 5x2 - 28x - 32 1 imaginary; 2 real 2 imaginary; 1 real 0 imaginary; 3 real 3 imaginary; 0 real

OpenStudy (anonymous):

Is the answer 3 imaginarys 0 reals

OpenStudy (anonymous):

if you factor this completely, you get (x+8)*(x-4)*(x+1) = 0

OpenStudy (anonymous):

this means that the zeros are -8 4 -1 so you have 3 real zeros, 0 imaginary zeros

OpenStudy (anonymous):

Now what about this; Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 8, -14, and 3 + 9i f(x) = x4 - 11x3 + 72x2 - 606x + 10,080 f(x) = x4 - 303x2 + 1212x - 10,080 f(x) = x4 - 11x3 - 72x2 + 606x - 10,080 f(x) = x4 - 58x2 + 1212x - 10,080

OpenStudy (acacia):

The rational roots theorem and some trivial analysis shows that -1 is a root. The first pair of coefficients and the last pair differ by the same amount (4). That means that (x+1) is a factor of the polynomial. Make that division. The result is a quadratic, to which you can apply the quadratic formula and finish the problem.

OpenStudy (anonymous):

still have no idea what that means